OpenStudy (anonymous):

hard question. help? (n+4)-3n=1+n

OpenStudy (anonymous):

(n+4)-3n=1+n -n -n (n+4)-4n = 1 solve for n

OpenStudy (dumbcow):

solve for n? this is basic algebra, not hard relatively speaking

OpenStudy (anonymous):

wait i dont get what you did

OpenStudy (anonymous):

@larryboxaplenty

OpenStudy (anonymous):

i can see you subtracted n from the left and right. now the n on the right canceled it self out and we have 1. but how did you subtract n from the left and end up with (n+4)-4n?